OpenStudy (anonymous):

1-4d is greater than or equal to 4-d Solve for d.

OpenStudy (agent0smith):

\[\large 1 - 4d \ge 4 -d\] you want to get d alone. Start by adding d to both sides (that'll get rid of -d on the left)

OpenStudy (anonymous):

\[1-5d \ge 4-2d\] ?

OpenStudy (agent0smith):

-d+d is zero -4d + d is -3d

OpenStudy (anonymous):

I'm confused.

OpenStudy (agent0smith):

-1+1 is zero... -4 + 1 is -3 same thing applies to -d+d and -4d + d

OpenStudy (anonymous):

So what does the equation look like now?

OpenStudy (agent0smith):

add d to both sides, and simplify this \[\large 1 - 4d +d \ge 4 -d + d\]

OpenStudy (anonymous):

So...what do I do now?

OpenStudy (agent0smith):

simplify it :P like i showed above... the -d+d is zero. -4d+d is -3d. So simplify the above

OpenStudy (agent0smith):

so the -d+d cancels off...what about the -4d+d?\[\large 1 - 4d +d \ge 4 \]

OpenStudy (anonymous):

1−4d+d≥4 1-4d+d-4≥_ ?

OpenStudy (agent0smith):

simplify the left, leave the 4 alone for now 1−3d≥4

OpenStudy (agent0smith):

now subtract 1 on both sides